OpenStudy (anonymous):

HELP PLEASE I already have the answer just need to double check. 15x-20y=7 10y=-3+5x

OpenStudy (anonymous):

x = 1/5 y= -1/5 should be it

OpenStudy (anonymous):

Do you understand how to get there?

OpenStudy (anonymous):

Yes I got that already I forgot to clear the question. thank you though!